An alloy of copper is 10% copper and weighs 25 pounds. A second alloy is 18% copper. How much (to the nearest tenth lb.) of the second alloy must be added to the first alloy to get a 13% mixture.
Answer by checkley77(12844)

.10*25+.18x=.13(25+x)
2.5+.18x=3.25+.13x
.18x-.13x=4.5-3.25
.05x=.75
x=.75/.05
x=15 pounds of 18% copper is needed.
Proof:
.10*25+.18*15=.13(25+15)
2.5+2.7=.13*40
5.2=5.2